Skip to main content

AAVSOnet news 2014-03-26

1 post / 0 new
HQA's picture
AAVSOnet news 2014-03-26

BSM_South is shut down for a week, as Peter Nelson is taking a holiday.

BSM_Hamren was originally installed and operated by Bob Stine in California.  Bob is selling his dark site, and so BSM_Hamren is being relocated to Hawaii.  Mike Linnolt has been kind enough to offer his assistance in getting this system going, and providing a unique longitude for bright-star observations.  I'll let you know when we get this system up and running, but it will be within the next couple of months.

BSM_Berry was temporarily installed on the roof of HQ.  Greg Bolt in Perth has offered to run this system, and so it will be packed and shipped during the next month.  Richard Berry has funded the purchase of additional filters for BSM_Berry, so it will have BVRI, griz and two SA200 gratings (oriented 90 degrees with respect to one another).   This system should also come on-line within a few months.

BSM_NM is up and running again.  The mount had gotten into a strange state that required Bill Stein's manual assistance to solve.

BSM HQ is a new system that is under configuration.  It will go on top of HQ where BSM_Berry lived, using a new mount and an AT65EDQ refractor.  This will be our test bed system, testing the SA200 as well as a possible single-board control system for all BSM sites.  It should be back on-line near the end of April.

OC61 has had some camera problems recently.  Our evaluation is that the camera is ok, but that some cabling is getting stretched in certain orientations of the mount.  We think that we've solved this problem, won't be sure until the system has consistently run several nights.

TMO61 has had tracking problems recently, where the telescope points correctly but exposures result in short trails.  Jon Holtzman is working on a solution.  We are sending Sloan g&i filters, along with an SA200, for installation in TMO61's filter wheel.

Coker30 is operating reasonably well, though it seems to have lost its pointing model recently.  Bill Goff has rebuilt that model and is waiting for clear weather to see if that solves the problem.

SRO50 is working reliably, though the Arizona weather has been particularly bad over the past couple of months.

Hardware updates for the network include new filters for BSM_Berry and TMO61; the inclusion of diffraction gratings on northern and southern telescopes, the new BSM_HQ setup, and work on the instrument selector for OC61.  We are planning to have the Eshel spectrograph in commissioning mode in a few months at Mt. John, followed with its twin in the northern hemisphere in the Fall.  We've had several new proposals covering dozens of targets that are being placed into various queues, and have started work on the epoch photometry databases for K28, K35, W28 and W30.


AAVSO 49 Bay State Rd. Cambridge, MA 02138 617-354-0484